A vessel containing an unknown volume of Argon in cm³at 300°F and 101kpa was inverted and placed in cold Ethanol. As the gas contracted , the pressure was quadrapled while the temperature was reduced by one _third.

What is the final volume of the gas in dm³


1
Expert's answer
2021-08-02T02:13:01-0400

V1=?

T1=300ºF   

P= 101kPa

 

V=?

T2=100°F

P2=404kPa


 PV = nRT

T1 = 300 0F = (5/9 x (300 - 32)) + 273 = 422 K;

T2 = 300/3 = 100 0F = 311 K

P1 = 101 kPa = 101000 Pa

P2 = 101 kPa x 4 = 404 kPa = 404000 Pa

P1V1 / T1 = P2V2 / T2

101000 x V1 / 422 = 404000 x V2 / 311

= 512.969882 m/s